Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 312     Entry time: 19 Aug 2023, 06:20
Author: Konopkova Zuzana 
Group: HED 
Subject: Shift summary 

beam parameters: 7.7 keV seeded, 2.2 MHz, up to 20 pulses, 150 uJ/pulse

  • user program, measuring IXS spectra of plasmons with high resolution using seeded beam and monochromators
  • mono (re-)alignment to minimize the SASE pedestal of the seed. Need to search for good position of the beam on the crystals as with just a slight change in pointing, the transmission goes down and deteriorates the beam spot
  • haveing troubles getting good transmission through the mono
  • taking spectra with the monochromatized seeded beam on PMMA, SiO2 and Al to be compared with the non-monochromatized cold spectra with 20 bunches.
  • second half of the train is often not lasing though transmission of the pulses through the warm mono seems ok
ELOG V3.1.4-7c3fd00